# Towards Generalist Robot Learning from Internet Video: A Survey

### v3

@@ Abstract
Scaling deep learning to huge internet-scraped datasets has yielded remarkably general capabilities in natural language processing and visual understanding and generation. In contrast, data is scarce and expensive to collect in robotics. This has seen robot learning struggle to match the generality of capabilities observed in other domains. Learning from Videos (LfV) methods seek to address this data bottleneck by augmenting traditional robot data with large internet-scraped video datasets. Such video data may provide the model with foundational information regarding physical behaviours and the physics of the world. This holds great promise for improving the generality of our robots.
In this survey, we present an overview of the emerging field of LfV. We outline fundamental concepts, including the benefits and challenges of LfV. We provide a comprehensive review of current methods for: extracting knowledge from large-scale internet video; tackling key LfV challenges; and boosting downstream reinforcement and robot learning via the use of video data. LfV datasets and benchmarks are also reviewed. The survey closes with a critical discussion of challenges and opportunities. Here, we advocate for scalable foundation model approaches that can leverage the full range of available internet video to aid the learning of robot policies and dynamics models. We hope this survey can inform and catalyse further LfV research, facilitating progress towards the development of general-purpose robots.
@@ 1 Introduction
Robot learning can take inspiration from advances in other domains. Deep learning has recently provided remarkable improvements in natural language processing (?), image generation (?), and video generation (?). This has been achieved by training expressive architectures (?) on massive, diverse datasets scraped from the internet. Here, performances have been shown to consistently and predictably improve with increased computing power and data (?). This use of diverse internet data has facilitated a transition from task-specific models to monolithic ‘foundation models’ with more general capabilities (?).
Scaling-up robot learning in a similar fashion is certainly a promising avenue towards obtaining more general-purpose robot capabilities (?, ?). However, obtaining large-scale data is a challenge in robotics. Robotics faces a chicken-and-egg problem: we cannot easily collect robot data due to limited robot capabilities (deploying these robots to collect real-world data can be ineffective and dangerous), and we cannot easily improve capabilities due to the lack of data (see Figure 1).
How can we overcome this critical data bottleneck? One possibility is to use human-teleoperation to collect robot data (?, ?). However, this is expensive and can be difficult in tasks requiring skill or dexterity. Another option is to leverage simulation (?, ?). However, simulation comes with several issues, including inaccurate simulated physics and difficulties regarding creating a suitable diversity of simulated environments and tasks. A final option is, like previous deep learning successes (?, ?), to leverage the vast quantities of data already available on the internet.
While any practical approach may leverage all of the above complementary sources of data, in this survey we focus on learning from internet data. Specifically, our interest lies in internet video data. Our reasoning here is threefold:
@@
    Relevant information content. The information content of internet video is highly relevant to a general-purpose robot. In contrast to text or image data, internet video can uniquely offer information regarding the physics and dynamics of the world, and information regarding human behaviours and actions (?). Crucially, internet video has excellent coverage over many relevant behaviours and tasks (e.g., household chores).
@@
By leveraging large, diverse internet video data during robot learning, we hope to obtain significantly better generalization beyond the robot data than approaches that rely solely on small, narrow robot datasets (see Figure 2 and Section 3.3). Indeed, recent progress in the emerging field of Learning from Videos (LfV) has been promising, demonstrating evidence of this. This has included work leveraging large-scale video prediction models to act as robot dynamics models (?, ?), and work leveraging both robot data and internet video to train foundational robot policies (?).
Nevertheless, leveraging internet video comes with a number of challenges (Section 3.4). First, video is a challenging data modality: it is high-dimensional, noisy, and poorly labelled. Second, video lacks information critical to robotics, including action labels and low-level force and proprioceptive information. Moreover, there may be various distribution shifts between internet video and the robot domain. Given these challenges, two key questions for LfV research are: ‘How to extract relevant knowledge from internet video?’ and ‘How to apply video-extracted knowledge to robotics?’.
In this survey, we review literature that attempts to answer these questions  [^1] See Figure 1 and the Table of Contents for more detail regarding the survey’s structure.. For the first question, we survey video foundation model techniques promising for extracting knowledge from large-scale internet video (Section 4.1). For the second, we perform a thorough analysis of literature that addresses LfV challenges (Section 4.2) and that utilises video data to aid robot learning (Section 4.3). We conclude the survey by discussing challenges and opportunities for future LfV research (Section 3.4). Here, we recommend future research to focus on scalable approaches that can leverage the full range of available internet video data, and to target the learning of policies and dynamics models data in order to best obtain the benefits of LfV.
Existing LfV research has shown promising signs of life, and this is a field ripe for further progress. While it is unclear whether scaling deep learning will take us all the way to general-purpose robots, it seems highly likely that scaling, with the help of internet video data, can provide significant gains in the near-future. We hope this comprehensive survey can encourage and inform future LfV research, serving to facilitate our progress towards the creation of generalist robots.
@@ 2 Background
Recent successes in deep learning (?, ?, ?, ?) have largely been driven by: (i) scaling dataset sizes; (ii) scaling model parameter sizes; (iii) improved model architectures – e.g., transformer (?) and diffusion models (?); (iv) and the use of self-supervised objective functions that allow for the use of large-scale, but unlabelled, internet data. Empirically-derived ‘scaling laws’ have shown deep learning performances to consistently and predictably improve with increased computation, model parameter count, and data (?). The term ‘foundation model’ is commonly used to refer to a large general-purpose model, pretrained on large diverse data, that can be used to solve wide varieties of downstream tasks (?).
Scaling deep learning has led to remarkable progress in natural language processing. Large language models (LLMs) are transformer models with billions of parameters, pretrained via a simple next-token prediction objective on huge textual datasets scraped from the internet (?, ?). This pretraining scheme allows the model to learn a ‘foundation’ of declarative and procedural knowledge. The pretrained LLM can be finetuned via supervised learning on high-quality curated data and/or via reinforcement learning from human (or AI) feedback (RLHF) (?, ?). This finetuning creates a more functional product for the end-user.
Similar trends can be seen in the domain of images and video. Image generation has benefited from the use of large-scale internet data and diffusion architectures (?, ?). Meanwhile, multi-modal vision-language models (VLMs), trained on large-scale data to take both images and text as input, have led to progress in image understanding (?, ?, ?). Note, we detail video foundation models efforts in Section 4.1.
@@
There is a hope that scaling robot learning may lead to successes similar to those seen in other domains, and thus improve the generality of our robots (?). However, there are well-known challenges here. Standard online reinforcement learning from scratch is not an easily scalable solution for real-world robotics: it is time-consuming, costly, and potentially dangerous (?). Offline learning (?, ?) offers a potential solution, but obtaining suitably large-scale and diverse datasets is non-trivial (?). In light of these issues, a number of proposed approaches have been proposed.
@@
  • Simulation. The use of simulation allows us to avoid the difficulties associated with real-world online learning (?, ?, ?). However, simulation presents a number of issues. (1) Inaccuracies in low-level simulation physics creates a ‘sim-to-real’ gap (?) that must be overcome. A partial solution here is to employ domain randomization (?). (2) Manually creating a suitable diversity of simulated environments and tasks for generalist robotics is a challenge. Recent works seek to tackle this using procedurally generated environments (?, ?) and foundation-model-assisted environment design (?, ?, ?). (3) We often lack a policy capable of collecting high-quality data in the simulated environment. Solutions here have included using humans to collect the data (?), or using LLMs with access to privileged simulation information as policies (?).
@@
  • Scaling real world data collection. Data collected via human teleoperation (?, ?) and pooled data from multiple academic labs (?) have been used to train robot foundation models (?, ?). Other works have investigated methods for automating data collection to improve scalablity (?, ?, ?). Several commercial companies have demonstrated evidence of infrastructure suitable for large-scale robot data-collection (?, ?). However, the largest open-source robot dataset (?) is still significantly smaller and less diverse than internet-scale data (see Figure 8).
@@
  • Internet data. Robot learning has benefited from the use of internet data. Image and video data have been used to pretrain visual representations for robotics (?, ?). Foundational VLMs and LLMs have been used to help define reward functions for the robot learner (?, ?, ?, ?). LLMs have been employed as high-level planners in long-horizons tasks (?, ?). Some works have finetuned a standard internet-pretrained foundation model into a robot foundation model (?, ?). Others have jointly pretrained on both internet data and action-labelled robot data (?, ?). We elaborate on how internet video data has been used to aid robot learning in Section 4.3.
@@
To the best of our knowledge, we are the first survey focusing explicitly on approaches that can scale-up general-purpose robot learning via the use of large quantities of diverse internet video data. ? (?, ?) survey imitating learning methods, ? (?) review of offline RL methods, while ? (?) highlight the promise of transferring pretrained knowledge from a source to a target domain in RL. ? (?, ?) review methods for utilising and developing foundation models for decision-making and robotic applications. ? (?) review self-supervised learning methods for video data, ? (?) review the use of foundation models for video understanding, while ? (?, ?) provide details on how LLMs can aid video understanding.
? (?) review imitation learning from observational data, while ? (?) advocate for the use of video (and video generation methods in particular) as a unified interface to absorb internet knowledge and represent diverse tasks. ? (?) present a survey of video-based learning approaches for robot manipulation. In contrast to this work, we place a stronger focus on approaches that: (i) can scale to large, diverse internet video data, and (ii) can yield general-purpose robotic capabilities.
@@ 3.2 Learning from Videos (LfV): Formalism and Assumptions
We assume an LfV method has access to a video dataset $D_{\text{video}}$. We denote a video clip as $\tau=(o_{0},o_{1},...,o_{T})$, where $\tau$ is the full clip and each $o$ is an RGB image observation. Optionally, $D_{\text{video}}$ may come paired with language annotations or annotations of other modalities. In general, we will also assume access to a robot dataset $D_{\text{robot}}$. This contains trajectories of transition tuples $(o_{t},a_{t},r_{t},o_{t+1})$, where $r_{t}$ may be missing, and where each $o$ contains an image observation and may also contain other information (e.g., tactile information). The goal of LfV is to leverage $D_{\text{video}}$ to improve robot performances versus when learning from $D_{\text{robot}}$ alone.
@@
Limitations of these assumptions. First, for certain robot tasks, embodiments very different from that of a human may prove more effective. Second, $D_{\text{video}}$ may not in fact have coverage over all the tasks we want the robot to perform: some important human actions and tasks may be underrepresented in the video dataset. Additionally, the robot may need to perform tasks not commonly performed by humans. Nevertheless, even in these cases internet video may still provide generally information regarding the world and physical behaviour.
@@ 3.3 Benefits of LfV
LfV offers the exciting possibility of generalizing beyond narrow robot datasets, $D_{\text{robot}}$, to the full space of tasks covered in $D_{\text{video}}$. The rationale behind this expectation is as follows. First, consider a $D_{\text{robot}}$ that contains most of the low-level skills or ‘atomic’ actions required from the robot (e.g., specific grasping motions or locomotion skills). Now consider a task unseen in $D_{\text{robot}}$ but seen in $D_{\text{video}}$. In this case, our combined dataset ($D_{\text{combined}}=D_{\text{robot}}+D_{\text{video}}$) contains most information required to complete the task. $D_{\text{robot}}$ provides information regarding how to execute the low-level skills, whilst $D_{\text{video}}$ provides higher-level information regarding how to complete the overall task (e.g., visual information regarding required movements and steps). Thus, a suitable LfV method may be capable of leveraging $D_{\text{video}}$ to generalise beyond $D_{\text{robot}}$ and solve the task. There is preliminary evidence of this in LfV-related literature (?, ?, ?, ?). Figure 2 explores this generalization setting in more detail.
@@ 3.4 Fundamental Challenges
We now discuss the key fundamental challenges that may be encountered when attempting to learn from internet video for robotics (see Figure 3 for visualizations). This provides context to understand the motivation behind specific LfV methods proposed in the literature (as we detail in Section 4).
@@
There may be a distributional shift between an internet video dataset and the downstream robot domain. This can include differences in physical embodiments, camera viewpoints, or environments. Additionally, humans in videos may perform behaviours which are sub-optimal or irrelevant to the downstream robot task. These shifts present a challenge to deep learning methods.

### v4

@@ Abstract
Scaling deep learning to massive, diverse internet data has yielded remarkably general capabilities in visual and natural language understanding and generation. However, data has remained scarce and challenging to collect in robotics, seeing robot learning struggle to obtain similarly general capabilities. Promising Learning from Videos (LfV) methods aim to address the robotics data bottleneck by augmenting traditional robot data with large-scale internet video data. This video data offers broad foundational information regarding physical behaviour and the underlying physics of the world, and thus can be highly informative for a generalist robot.
In this survey, we present a thorough overview of the emerging field of LfV. We outline fundamental concepts, including the benefits and challenges of LfV. We provide a comprehensive review of current methods for extracting knowledge from large-scale internet video, addressing key challenges in LfV, and boosting downstream robot and reinforcement learning via the use of video data. The survey concludes with a critical discussion of challenges and opportunities in LfV. Here, we advocate for scalable foundation model approaches that can leverage the full range of available internet video to improve the learning of robot policies and dynamics models. We hope this survey can inform and catalyse further LfV research, driving progress towards the development of general-purpose robots.
@@ 1 Introduction
Robot learning can take inspiration from advances in other domains. Deep learning has recently provided remarkable improvements in natural language processing (?), image generation (?), and video generation (?). This has been achieved by training expressive architectures (?) on massive, diverse datasets scraped from the internet. Here, ‘scaling laws’ have shown performances to consistently and predictably improve with increased computational power and data (?). The use of diverse internet data has facilitated a transition from task-specific models to monolithic ‘foundation models’ with more general capabilities (?).
There is evidence that these deep learning techniques (?, ?) and scaling laws (?, ?, ?) can transfer to robotics and control. This offers a path towards more general-purpose robot capabilities. However, obtaining large-scale data is a challenge in robotics. Robotics faces a chicken-and-egg problem: data cannot be easily collected due to limited robot capabilities (deploying limited robots to collect real-world data can be ineffective and dangerous), and capabilities cannot easily be improved due to the lack of data (see Figure 1).
How can we overcome this data bottleneck? One possibility is to use humans to collect robot data (?, ?). However, this is expensive and can be difficult in tasks requiring skill. Another option is to leverage simulation (?, ?). However, simulation comes with issues related to flawed simulated physics and difficulties creating and training on a suitable diversity of simulated environments and tasks. A final option is, like previous deep learning successes (?, ?), to leverage the vast quantities of data available on the internet.
While any practical approach may leverage all complementary sources of data, in this survey we focus on learning from internet data. Specifically, our interest lies in internet video data. Our reasoning here is threefold:
@@
    Relevant information content. In contrast to internet-scraped text or image data, internet video can uniquely offer information regarding the physics and dynamics of the world, and information regarding human behaviours and actions (?). Crucially, internet video has coverage over many behaviours and tasks relevant to a general-purpose robot (e.g., household chores).
@@
By leveraging large, diverse internet video data during robot learning, we hope to obtain a number of benefits (Section 3.3). This includes obtaining improved generalization beyond the available robot data, versus approaches that rely solely on small, narrow robot datasets (see Figure 2). Indeed, recent progress in the emerging field of Learning from Videos (LfV) (Section 3.2) has been promising, demonstrating evidence of this. This has included work leveraging large-scale video prediction models to act as robot dynamics models (?, ?), and work leveraging both robot data and internet video to train foundation models for robotics (?).
Nevertheless, leveraging internet video is non-trivial and comes with a number of challenges (Section 3.4). First, video is a challenging data modality: it is high-dimensional, noisy, and poorly labelled. Second, video lacks information critical to robotics, including action labels and low-level force and proprioceptive information. Moreover, there may be various distribution shifts between internet video and the robot domain. Given these challenges, two key questions for LfV research are: (i) ‘How to extract relevant knowledge from internet video?’ and (ii) ‘How to apply video-extracted knowledge to robotics?’.
In this survey, we review literature that attempts to answer these questions  [^1] See Figure 1 and the Table of Contents for more details regarding the structure of the survey.. For the first question, we survey video foundation model techniques promising for extracting knowledge from large-scale, heterogenous internet video (Section 4.1). These techniques may serve as the backbone of scalable LfV approaches. For the second question, we perform a thorough analysis of literature that explicitly addresses LfV challenges (Section 4.2), before reviewing methods that directly utilise video data to aid robot learning (Section 4.3). We conclude the survey by discussing challenges and opportunities for future LfV research (Section 6). Here, we recommend focusing on scalable approaches that can leverage the full range of available internet video data, and focusing on learning policies and dynamics models to best obtain the benefits of LfV.
Existing LfV research has shown promising signs of life, and this is a field ripe for further progress. Of course, it is unclear whether scaling deep learning will take us all the way to general-purpose robots (Section 6.4), and there are other research directions that should be pursued in parallel (Section 2). Nevertheless, with sufficient effort and innovation, it seems likely that scaling—with the help of internet video data—can provide significant gains in the near-future. We hope this comprehensive survey can encourage and inform these future LfV research efforts, serving to facilitate our progress towards the creation of generalist robots.
@@ 2 Background
Empirically-derived ‘scaling laws’ have shown deep learning performances to consistently and predictably improve with increased computation, model parameter count, and data (?). Recent successes in deep learning (?, ?, ?, ?) have largely been driven by such scaling, aided by the use of improved model architectures – transformer (?) and diffusion models (?), and the use of self-supervised objectives that allow for learning from unlabelled internet data. The term ‘foundation model’ is commonly used to refer to a large, general-purpose model, pretrained on large diverse data, that can solve wide varieties of downstream tasks (?).
Large language models (LLMs) are transformer models with billions of parameters (?), pretrained via a simple next-token prediction objective on huge text datasets scraped from the internet (?, ?). This pretraining scheme allows the model to learn a ‘foundation’ of declarative and procedural knowledge. The pretrained LLM can be finetuned into a more functional product via supervised learning on high-quality curated data and/or via reinforcement learning from human (or AI) feedback (RLHF) (?, ?). Meanwhile, image generation has also benefited from the use of large-scale internet data and from diffusion architectures (?, ?). Multi-modal vision-language models (VLMs), trained on large-scale internet data to take both images and text as input, have progressed visual understanding (?, ?, ?). We detail video foundation model efforts in Section 4.1.
@@
There is a hope that scaling robot learning, in a similar fashion to other domains, may improve the generality of our robots (?). However, there are well-know challenges here. Unlike in video games (?, ?), scaling online reinforcement learning (RL) is not a practical solution for real-world robotics: it is time-consuming, costly, and dangerous (?). Meanwhile, offline learning solutions (?, ?) lack suitably large-scale, high-quality robotic datasets (?). A number of approaches have been proposed to address these challenges and scale up robotic datasets:
@@
  • Simulated data. The use of simulation is promising for scaling up online learning while avoiding avoiding difficulties associated with the real-world (?, ?, ?). However, the use of simulation presents a number of issues (?). (1) Inaccuracies in low-level physics create a ‘sim-to-real’ gap (?) that must be overcome. A partial solution here is to employ domain randomization (?). (2) Manually creating a suitable diversity of simulated environments and tasks for generalist robotics is a challenge. Recent works seek to tackle this using procedural environment generation (?, ?) and foundation-model-assisted environment design (?, ?, ?). (3) We often lack a policy to collect high-quality data in the simulation. Solutions here have included the use of an automated curriculum during RL training (?, ?) and collecting data using: specialist policies (?), models with access to privileged simulation information (?, ?, ?), or bootstrapping data collection from human demonstrations (?, ?).
@@
  • Real world data. Data collected via human teleoperation (?, ?) and pooled data from multiple academic labs (?) have been used to train initial robot foundation models (?, ?). Other works have investigated methods for automating data collection to improve scalablity (?, ?, ?). Several companies have demonstrated evidence of infrastructure suitable for large-scale robot data-collection (?, ?). However, the largest open-source robot dataset (?) is still significantly smaller and less diverse than internet-scale data (see Figure 8).
@@
  • Internet data. Robot learning has benefited from the use of broad internet data. Image and video data have been used to pretrain visual representations for robotics (?, ?). Foundational VLMs and LLMs have been used to help define reward functions for the robot learner (?, ?, ?, ?). LLMs have been employed as high-level planners in long-horizons tasks (?, ?). Some works have finetuned a standard internet-pretrained foundation model into a robot foundation model (?, ?). Others have jointly pretrained on both internet data and action-labelled robot data (?, ?). We elaborate on how internet video data has been used to aid robot learning in Section 4.3.
Other Paths to Generalist Robots.
Scaling end-to-end deep learning approaches that train a monolithic ‘foundation’ model is one potential path to general-purpose robot capabilities. Other paths could involve combining learning with the use of stronger inductive biases, more structure, or classical robotics techniques. These may help address robotics-specific challenges and potential limitations of scaling end-to-end deep learning (?, Section 6.3). Modular approaches have been proposed which use specialist models for different skills or tasks (?, ?), or use a hierarchy to separate out high-level planning from low-level control (?, ?, ?). Learning has been combined with explicit planning algorithms (?, ?) and environment representation techniques (?, ?, ?). Structure related to the robot embodiment (?, ?) or the physics of the system (?, ?) can be leveraged, while some approaches carefully choose and process their observation and action spaces (?, ?). As these ‘alternative’ approaches and deep learning are often combined, they can be complementary: advances in scaling deep learning can benefit alternative approaches, and visa versa (e.g., alternative approaches can collect data for monolithic deep learning models; ?).
@@
? (?, ?) survey imitation learning methods; ? (?) review offline RL methods; while ? (?) highlight the promise of transferring knowledge from a source to a target domain in RL. ? (?, ?) review the use of foundation models in decision-making and robotics. Relevant works in the video ML literature have included reviews of self-supervised video learning (?), foundation models for video understanding (?), and LLMs for video understanding (?, ?).
? (?) review imitation learning from observational data, while ? (?) advocate for the use of video (and video generation) as a unified interface to absorb internet knowledge and represent diverse tasks. ? (?) review video-based learning approaches for robot manipulation. In contrast to this work, our survey places a stronger focus on approaches that: (i) can scale to large, diverse internet video data, and (ii) can yield general-purpose robotic capabilities.
@@ 3.1 Reinforcement Learning (RL)
In the LfV setting, the POMDP is some real-world environment and $\pi(a_{t}|s_{t})$ controls a robot within this environment.
@@ 3.2 Learning from Videos (LfV): Formalism and Assumptions
We assume an LfV method has access to a video dataset $D_{\text{video}}$. We denote a video clip as $\tau=(o_{0},o_{1},...,o_{T})$, where $\tau$ is the full clip and each $o$ is an RGB image observation. Optionally, $D_{\text{video}}$ may come paired with language annotations or annotations of other modalities. We also assume access to a robot dataset $D_{\text{robot}}$. This contains trajectories of transition tuples $(o_{t},a_{t},r_{t},o_{t+1})$, where $r_{t}$ may be missing, and where each $o$ contains an image observation and may also contain other information (e.g., tactile information). The goal of LfV is to leverage our combined LfV dataset, $D_{\text{LfV}}=\{D_{\text{robot}},D_{\text{video}}\}$, to obtain an improved $\pi(a_{t}|s_{t})$ versus when learning from $D_{\text{robot}}$ alone.
@@
Limitations of these assumptions. First, for certain robot tasks, embodiments very different from that of a human may prove more effective. Second, $D_{\text{video}}$ may not in fact have coverage over all the tasks we want the robot to perform: some important actions and tasks may be underrepresented in the video dataset. Additionally, the robot may need to perform tasks not commonly performed by humans. Nevertheless, even in these cases, internet video can still provide generally information regarding the world and physical behaviour.
@@ 3.3 Benefits of LfV
LfV offers the exciting possibility of generalizing beyond narrow robot datasets, $D_{\text{robot}}$, to the full space of tasks covered in the video dataset, $D_{\text{video}}$. The rationale behind this expectation is as follows. First, consider a $D_{\text{robot}}$ that contains most of the low-level skills or ‘atomic’ actions required from the robot (e.g., specific grasping motions or locomotion skills). Now consider a task unseen in $D_{\text{robot}}$ but seen in $D_{\text{video}}$. In this case, our combined LfV dataset ($D_{\text{LfV}}=\{D_{\text{robot}},D_{\text{video}}\}$) contains most information required to complete the task. $D_{\text{robot}}$ provides information regarding how to execute the low-level skills, whilst $D_{\text{video}}$ provides higher-level information regarding how to complete the overall task (e.g., visual information regarding required movements and steps). Thus, a suitable LfV method may be capable of leveraging $D_{\text{video}}$ to generalise beyond $D_{\text{robot}}$ and solve the task. There is preliminary evidence of this in LfV-related literature (?, ?, ?, ?). Figure 2 explores this generalization setting in more detail.
@@ 3.4 Fundamental Challenges
Learning from internet video for robotics comes with some fundamental challenges. These render LfV a non-trivial problem. Awareness of these challenges is crucial for understanding the motivation behind existing LfV research (Section 4) and the future promising research directions we highlight (Section 6). Note, none of these challenges have yet been fully solved.
@@
There may be significant distributional shifts between an internet video dataset and the downstream robot domain. This can include differences in physical embodiments, camera viewpoints, tasks, and environments. Humans in videos may perform behaviours which are sub-optimal or irrelevant to the downstream robot. Certain useful tasks and behaviours may be underrepresented in internet video (e.g., coverage over specific factory tasks may be poor). These shifts present a challenge to deep learning methods.